OpenStudy (kaylala):

Henry counted the money he saved for one month and found he has $8990 consisting of $100, $50 and $20 bills. He had half as many $50 bills as $20 bills but2 times as many as $100 bills as $50 bills. How many bills of each kind did he have?

OpenStudy (kaylala):

Please teach me how. Thanks

OpenStudy (anonymous):

let $20 bills are x then can you tell how many bills of $50 he had?

OpenStudy (kaylala):

.5x = 50 bills 2(.5x)= 100 bills Is that right? Then how do you solve it?

OpenStudy (kaylala):

@03225186213 ^

OpenStudy (anonymous):

no he has half of as many $50 bills as $20 so $ 50 bills are x/2 can you tell how many $100 bills he has?

OpenStudy (kaylala):

But isnt .5 = 1/2 Well okay then 100 bills =2(x/2)

OpenStudy (anonymous):

you are right i did not look dot in it

OpenStudy (kaylala):

Okay

OpenStudy (anonymous):

now make equation with this information.

OpenStudy (kaylala):

How? x+.5x+2(.5x) = ?

OpenStudy (kaylala):

@03225186213 ^

OpenStudy (anonymous):

no 20x+(.5x)(50)+100x=8990 now solve for x

OpenStudy (kaylala):

X=62 Then what @03225186213

OpenStudy (anonymous):

henry has 62 bills of $20 31 bills of $50 and 62 bills if $100